# Unlocking Video-LLM via Agent-of-Thoughts Distillation

### v1

@@ 1 Introduction
Video Question Answering (VideoQA) refers to a critical task that offers a natural interface for human-machine interaction through language [47, 41, 42, 33]. This synergy of visual content and language enhances the accessibility of AI systems for the general public, allowing users to query complex visual content with everyday language. By encompassing tasks such as action recognition, object detection, and scene understanding, VideoQA serves as a comprehensive benchmark for evaluating AI’s ability to interpret videos, addressing the fundamental questions of ‘who’, ‘what’, ‘when’, and ‘where’ that are crucial to understand daily life activities, pushing the boundaries of what AI systems can interpret from dynamic visual content.
@@ 3.4 Step-by-step Distillation
Model MVBench VideoMME STAR NExT-QA Perception-Test
(Acc.) (Acc.) (Acc.) (Acc.) (Acc.)
@@
LLaMA-VID (7B) [21] 41.9 25.9 - - 44.6
Video-LLaVA (7B) [22] 41.0 39.9 - - 44.3
VideoChat2 (7B) [19] 51.1 33.7 59.0* 68.6* 47.3
VideoLLaMA2 (7B) [4] 53.4 45.1 58.5* 62.3* 49.6
LLaVA-NeXT-Video (7B) [52] 46.5* 41.0* 52.4* 61.6* 47.5*
LLaVA-NeXT-Video-Instruct (7B) 53.4 43.2 72.2 77.1 50.3
LLaVA-NeXT-Video-AoTD (7B) 55.6 45.0 74.3 77.6 50.6
Table 4: Comparison with Video-LLMs on MC-VQA benchmarks. LLaVA-NeXT-Video-AoTD outperforms all other baselines the and the version without CoT distillation. * means results reproduced by ourseleves. Results without signs are retrieved from [4].
@@ 4.1 Experimental Setup
Evaluation benchmarks. We conduct extensive evaluations on Multiple-Choice Video QA (MC-VQA) and Open-Ended Video QA (OE-VQA). We report the top-1 accuracy for all MC benchmarks, which means the proportion of the output equal to the answer. We report a GPT-assessed Acc. and Score with the help of GPT-3.5-turbo-0613 for all OE benchmarks. For each question, GPT delivers a binary decision indicating whether the output is correct or incorrect, along with a similarity score reflecting the degree of alignment between the output and the correct answer. The term ‘Acc.’ refers to the percentage of correct outputs, while ‘Score’ represents the average similarity scores. For the evaluation on AGQA and ANetQA, due to the large volume of test set, we test on a subset of samples. We evenly select the benchmark in-domain and out-of-domain for testing to ensure a comprehensive and reasonable evaluation of the model capability. Detailed statistics for evaluation benchmarks are shown in Table 3.
@@ 4.2 Quantitative Results
We divide the comparison into two parts: the first focuses on comparing the distilled model with other baselines, while the second examines the difference between the instruct version and the AoTD version. Note that, the latter part will be mainly compared and discussed, to demonstrate the model’s improvement relative to its previous performance, as well as establishing the transferability of the method across different models.
@@ 4.3 Ablation Study
Analysis on model transferability. As AoTD is a distillation method that leverages Chain-of-Thoughts (CoTs), it can theoretically be applied to any Video-LLMs. To assess the transferability of our method, we conduct experiments on another very recent model, LLaVA-OneVision (7B) [3]. As shown in Table 6, our method also demonstrates significant improvements on the benchmarks, showing the transferability and robustness of the approach. Due to the rapid advancements in the computer vision field, evaluating all models and benchmarks is prohibitively infeasible. Thus, we focus on assessing a single model against selected benchmarks to provide a representative evaluation.

### v2

# Enhancing Video-LLM Reasoning via Agent-of-Thoughts Distillation

@@ 1 Introduction
Video Question Answering (VideoQA) refers to a critical task that offers a natural interface for human-machine interaction through language [48, 41, 42, 33]. This synergy of visual content and language enhances the accessibility of AI systems for the general public, allowing users to query complex visual content with natural language. By encompassing tasks such as action recognition, object detection, and scene understanding, VideoQA serves as a comprehensive benchmark for evaluating AI’s ability to interpret videos, addressing the fundamental questions of ‘who’, ‘what’, ‘when’, and ‘where’ that are crucial to understand daily life activities, pushing the boundaries of what AI systems can interpret from dynamic visual content.
@@ 3.4 Step-by-step Distillation
Model MVBench VideoMME STAR NExT-QA Perception-Test VSIBench
(Acc.) (Acc.) (Acc.) (Acc.) (Acc.) (Acc.)
@@
LLaMA-VID (7B) [21] 41.9 25.9 - - 44.6 -
Video-LLaVA (7B) [22] 41.0 39.9 - - 44.3 -
VideoChat2 (7B) [19] 51.1 33.7 59.0* 68.6* 47.3 -
VideoLLaMA2 (7B) [4] 53.4 45.1 58.5* 62.3* 49.6 -
LLaVA-NeXT-Video (7B) [53] 46.5* 41.0* 52.4* 61.6* 47.5* 19.7*
LLaVA-NeXT-Video-Instruct (7B) 53.4 43.2 72.2 77.1 50.3 26.7
LLaVA-NeXT-Video-AoTD (7B) 55.6 45.0 74.3 77.6 50.6 28.8
Table 4: Comparison with Video-LLMs on MC-VQA benchmarks. LLaVA-NeXT-Video-AoTD outperforms all other baselines the and the version without CoT distillation. * means results reproduced by ourseleves. Results without signs are retrieved from [4] and [43].
@@ 4.1 Experimental Setup
Evaluation benchmarks. We conduct extensive evaluations on Multiple-Choice Video QA (MC-VQA) and Open-Ended Video QA (OE-VQA). We report the top-1 accuracy for all MC benchmarks, which means the proportion of the output equal to the answer. We report a GPT-assessed Acc. and Score with the help of GPT-3.5-turbo-0613 for all OE benchmarks. For each question, GPT delivers a binary decision indicating whether the output is correct or incorrect, along with a similarity score reflecting the degree of alignment between the output and the correct answer. The term ‘Acc.’ refers to the percentage of correct outputs, while ‘Score’ represents the average similarity scores. For the evaluation on AGQA and ANetQA, due to the large volume of test set, we test on a subset of samples. We evenly select the benchmark in-domain and out-of-domain for testing to ensure a comprehensive and reasonable evaluation of the model capability. Noted that though VSIBench has both MC and OE questions, it doesn’t need GPT for score, so we classify it into MC benchmarks for convenience. Detailed statistics for evaluation benchmarks are shown in Table 3.
@@ 4.2 Quantitative Results
We divide the comparison into two parts: the first focuses on comparing the distilled model with other baselines, while the second examines the difference between the instruct version and the AoTD version. Note that, the latter part will be mainly compared and discussed, to demonstrate the model’s improvement relative to its previous performance, as well as establishing the transferability of the method across models.
@@ 4.3 Ablation Study
Analysis on model transferability. As AoTD is a distillation method that leverages Chain-of-Thoughts (CoTs), it can theoretically be applied to any Video-LLMs. To assess the transferability of our method, we conduct experiments on another very recent model, LLaVA-OneVision (7B) [3]. As shown in Table 7, our method also demonstrates significant improvements on the benchmarks, showing the transferability and robustness of the approach. Due to the rapid advancements in the computer vision field, evaluating all models and benchmarks is prohibitively infeasible. Thus, we focus on assessing some representative models against selected benchmarks to provide a representative evaluation.
Analysis on latency and computation. We compare the agent-based system with the distilled model (LNV-AoTD) across three key metrics. As shown in Table 6, LNV-AoTD outperforms the agent-based system in both inference latency and computational efficiency, demonstrating the necessity of distillation process.
Model Time (s) $\downarrow$ Memory (GB) $\downarrow$ TFLOPs $\downarrow$
Agent-based system 47.93 65.34 30.98
LNV-AoTD 10.58 18.21 13.53
Table 6: Latency and computation comparison.
@@
VideoLLaMA2-Instruct - 54.9 69.2 56.0/3.5
VideoLLaMA2-AoTD 56.0 71.1 57.2/3.5
Qwen2-VL-Instruct - 65.6 71.4 59.8/3.6
Qwen2-VL-AoTD 66.5 73.1 61.2/3.7